数理解析方法論 2 単位 (選択)

Methods for analysis of mathematical phenomena  今井仁司 · 教授/ 知的力学システム工学専攻建設創造システム工学コース社会システム工学講座

金成海 · 教授/ 国際センター

【授業目的】数理現象を解析するための手法を学ぶ．【授業概要】数理現象を解析するための手段として数値解析は欠かせない．数値

解析に用いられている数学的手法を講義する．工業にかかわる科目である．【授業形式】講義【キーワード】数理，数値解析【先行科目】 [先行科目] 【関連科目】『計算数理特論』 (0.5) 【履修要件】「基礎数学」の履修を前提とする．【履修上の注意】授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時

間の復習をしたうえで授業を受けることが，授業の理解と単位取得のために必要である．

【到達目標】 2 次元ポワソン方程式の境界値問題の適切性が差分法等を用いて理

解できる．【授業計画】

1. 計算機概論

2. 数値計算の常識

3. 高速計算法 (並列計算)

4. 差分法 I

5. 差分法 II

6. 差分法 III

7. 差分法 IV

8. 有限要素法 I

9. 有限要素法 II

10. 有限要素法 III

11. 有限要素法 IV

12. 境界要素法 I

13. 境界要素法 II

14. 境界要素法 III

15. 反復法【成績評価基準】授業中に課すレポートで評価する．【教科書】 [教科書] 【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216715 【対象学生】 [対象学生] 【連絡先】

⇒ 今井(A棟 220， 088-656-7541，携帯電話やE-mail での問い合わせは受け付けない) (オフィスアワー: オフィスアワー:木曜 14:00∼ 15:00)

Methods for analysis of mathematical phenomena 2 units (selection)

Hitoshi Imai · Professor / Planning and Design Systems Engineering for Infrastructures, Civil and Environmental Engineering, Intelligent Structures and Mechanics Systems Engineering

, Cheng-Hai Jin · Professor / International Center

Target⟩ To learn the to analyze the mathematical phenomena. Outline⟩ Methods used in analysis of mathematical phenomena are introduced.

Especially, those in numerical analysis are focused on. Style⟩ Lecture �������⟩ mathematics, numerical analysis Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “Advanced Computational Science”(0.5) Requirement⟩ Only the premise that have studied basic mathematics. Notice⟩ 授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復

習をしたうえで授業を受けることが，授業の理解と単位取得のために必要である．

Goal⟩ The adequacy of the two-dimensional boundary value problem governed by

the Poisson equation can be understood by the finite difference method, etc.

Schedule⟩

1. Introduction to computer

2. Common sense in numerical computation

3. High-speed computation (Parallel computing)

4. Finite difference method I

5. Finite difference method II

6. Finite difference method III

7. Finite difference method IV

8. Finite element method I

9. Finite element method II

10. Finite element method III

11. Finite element method IV

12. Boundary element method I

13. Boundary element method II

14. Boundary element method III

15. Iterative method Evaluation Criteria⟩ Evaluation by the report. Textbook⟩ [教科書] Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216715 Student⟩ [対象学生] Contact⟩

⇒ Imai(A220, +81-88-656-7541, The inquiry by means of the cellular phone or E-mail is not acceptable) (Office Hour: Office hours: Thursday 14:00-1 5:00)